# Cops and Robbers Pebbling in Graphs

### v3

@@ 1 Introduction
We define these three paradigms more specifically in Subsection 1.1 below. The new graph invariant we define to study in this paper is the cop pebbling number of a graph, denoted ${{\pi}^{\sf c}}(G)$; roughly, this equals the minimum number of pebble-cops necessary to capture the robber in the Cops and Robbers Pebbling paradigm. In Subsections 1.21.4 we present known results about Cops and Robbers, dominating sets, and optimal pebbling, respectively, that will be used in the sequel. We record in Subsections 2.12.3 new theorems on lower bounds, upper bounds, and exact answers for ${{\pi}^{\sf c}}(G)$, respectively, for a range of graph families, including paths, cycles, trees, chordal graphs, high girth graphs, and cop-win graphs, as well as, in some cases, for all graphs. For example, Theorem 14 proves that the cop pebbling number of $n$-vertex trees is at most $\lceil 2n/3\rceil$, which is tight for paths (and cycles), and Theorem 15 provides an upper bound involving a domination parameter, as a function of girth. Section 3 contains theorems for Cartesian products of graphs. For example, Theorems 22 and 23 give upper and lower bounds on the cop pebbling numbers of grids and cubes, respectively, while Theorem 24 proves that ${{\pi}^{\sf c}}(G\Box K_{t})\leq t{{\pi}^{\sf c}}(G)$ for all $G$. Furthermore, Theorems 26 and 27 show that the analogous inequality from Graham’s Pebbling Conjecture fails for cop pebbling. We finish in Section 4 with some natural questions left open from this work, including a version of Meyniel’s Cops and Robbers Conjecture that may hold in the Cops and Robbers Pebbling world, namely that ${{\pi}^{\sf c}}(G)\leq 2n/3+o(n)$ for all $G$.
@@ 1.1.3 Cop Pebbling
The cop pebbling number ${{\pi}^{\sf c}}(G)$ is defined as the minimum number $m$ so that, from some configuration of $m$ cops, it is possible to capture any robber via pebbling steps. For example, $c(G)\leq n(G)$, since the placement of one cop on each vertex has already caught any robber. We call an instance of a graph $G$, configuration $C$, and robber vertex $v$ a game, and say that the cops win the game if they can capture the robber; else the robber wins. Note that, since we lose a cop with each pebbling step, the cops must catch the robber within at most $|C|-1$ turns to win the game.
@@ 1.2 Cop Results
Here we list the results on Cops and Robbers that will be used to prove our theorems on cop pebbling. A graph $G$ is cop-win if $c(G)=1$. A vertex $u$ in $G$ is called a corner if there is a vertex $v\not=u$ such that $N[u]\subseteq N[v]$. We say that $G$ is dismantlable if either $G$ is a single vertex or there is a corner $u$ such that $G-u$ is dismantlable. Note that chordal graphs are dismantlable.
@@
[1] For $t\geq 1$ let $G$ be a graph with ${\sf gir}(G)\geq 8t-3$. Then $c(G)>({\delta}(G)-1)^{t}$.
Result 3.
[8] For $s\geq 1$ let $G$ be a graph with ${\sf gir}(G)\geq 4s+1$. Then $c(G)\geq\frac{1}{se}({\delta}(G)-1)^{s}$.
The latter two results are useless when ${\delta}(G)\leq 2$, so we use them only when ${\delta}(G)\geq 3$. Then, for $s=t=1$, Result 2 is stronger than Result 3. More generally, for $s=2t-1$, Result 3 is stronger than Result 2 whenever $({\delta}(G)-1)^{t-1}>(2t-1)e$. This holds when ${\delta}(G)=3$ and $t\geq 6$, when ${\delta}(G)=4$ and $t\geq 4$, when ${\delta}(G)\geq 5$ and $t\geq 3$, and when ${\delta}(G)\geq 10$ and $t\geq 2$. Additionally, when $s=2t$, Result 3 yields the lower bound $\frac{1}{2te}({\delta}(G)-1)^{2t}$, compared to the $({\delta}(G)-1)^{t}$ bound from Result 2, and so is stronger whenever $({\delta}(G)-1)^{t}>2te$. This holds when ${\delta}(G)=3$ and $t\geq 5$, when ${\delta}(G)=4$ and $t\geq 4$, and when ${\delta}(G)\geq 5$ and $t\geq 2$.
Result 4.
@@
Result 5 yields the following result as a corollary.
Result 6.
@@
Result 6 for even $d=2k$ also follows from the following result because $Q^{2}=C_{4}$.
Result 8.
[23] If $G=\Box_{i=1}^{k}C_{n_{i}}$ and each $n_{i}\geq 4$ then $c(G)=k+1$. [7] If $n_{i}=m$ for all $i$, then ${\sf capt}_{2k}(G)\leq k\lfloor\frac{m-1}{2}\rfloor(\lceil\textup{lg}\hskip 2.84526ptk\rceil+1)+1-k$.
In fact, Result 6 follows as well from the following result because $Q^{1}=P_{2}$ is a tree.
Result 9.
If each $T_{i}$ is a tree on at least two vertices and $G=\Box_{i=1}^{d}T_{i}$ then [21] $c=c(G)=\lceil\frac{d+1}{2}\rceil$ and [7] ${\sf capt}_{c}(G)\leq\lceil\lg d\rceil\sum_{i=1}^{d}{\sf rad}(T_{i})-\lfloor\frac{d-1}{2}\rfloor+1$.
@@ 1.3 Dominating Set Results
Result 10 implies that ${\gamma}(Q^{d})\geq 2^{d}/(d+1)$. The following result shows that the actual value is not much greater, asymptotically.
Result 13.
@@ 1.4 Optimal Pebbling Results
Result 17.
[9] If $G$ is an $n$-vertex graph with ${\sf gir}(G)\geq 2s+1$ and ${\delta}(G)=k$ then ${{\pi}^{*}}(G)\leq\frac{2^{2s}n}{{\sigma}_{k}(s)}$, where ${\sigma}_{k}(s)=1+k\sum_{i=1}^{s}(k-1)^{i-1}$.
@@
The next three results involve Cartesian products.
Result 20.
[17] For all $3\leq k\leq m$ we have ${{\pi}^{*}}(P_{k}\Box P_{m})\leq\frac{2}{7}km+8\approx 0.2857km$.
The above result is conjectured to be best possible, while the best known lower bound is below.
Result 21.
@@
Combining Results 15 and 16 yields the lower bound of the following result. The upper bound places $2^{k}$ pebbles on each vertex of a distance-$k$ dominating set, with $k$ roughly $d/3$.
Result 22.
@@ 2.1 Lower Bounds
If ${{\pi}^{\sf c}}(G)=c(G)$ then capturing the robber requires no steps. That is, if a pebbling capture requires a pebbling step, then the cop lost during the step is irrelevant in the Cops and Robbers capture, which contradicts the equality. That means that a successful pebbling configuration has no vertex without a pebble; i.e. ${{\pi}^{\sf c}}(G)=n(G)$. However, if $n(G)\geq 3$ then, by placing two cops at a vertex $v$ of degree at least two and one cop on each vertex not adjacent to $v$, then we can capture any robber in one step; thus ${{\pi}^{\sf c}}(G)<n$, a contradiction. (This is recorded in Corollary 6, below.) If $n(G)=2$ then $G=K_{2}$ and ${{\pi}^{\sf c}}(K_{2})=2>1=c(K_{2})$, a contradiction. Hence $G=K_{1}$. ∎
@@
For any graph $G$, ${{\pi}^{\sf c}}(G)\geq{{\pi}^{*}}(G)$. Equality holds if $G$ is a tree or cycle.
@@
Any configuration of $k$ cops, where $k<{{\pi}^{*}}(G)$, will contain a vertex $v$ which is unreachable. The robber can then choose to start and stay on $v$ and thus not be captured.
If $G$ is a cycle then Result 14 and Theorem 18 yield the equality.
If $G$ is a tree then place ${{\pi}^{*}}(G)$ cops according to an optimal pebbling configuration $C$. The robber beginning at some vertex $v$ defines a subtree $T$ containing $v$ for which every cop in $T$ is on a leaf of $T$, and any leaf of $T$ with no cop is a leaf of $G$. Thus the robber can never escape $T$. Because $C$ can reach every vertex of $T$, they can capture the robber, regardless of where he moves. Hence ${{\pi}^{\sf c}}(G)\leq{{\pi}^{*}}(G)$, and the equality follows. ∎
@@
If $G$ is a graph with ${\delta}={\delta}(G)\geq 27$, and with ${\sf gir}(G)\geq 4t+1$ and $n(G)\leq{\delta}^{2t+1}$ for some $t\geq 3$, then $c(G)>{{\pi}^{*}}(G)$.
@@
Given $G$ as above, set $d={\delta}-1$. Then we have ${\sf gir}(G)\geq 4t+1$, and so $c(G)\geq d^{t}/et$ by Result 3. Also, with $s=2t$, we have ${\sf gir}(G)\geq 2s+1$, and so ${{\pi}^{*}}(G)\leq 2^{2s}n/{\sigma}_{\delta}(s)$ by Result 17. Note that $2^{2s}n/{\sigma}_{\delta}(s)<4^{s}n/d^{s}\leq(4/d)^{2t}(d+1)^{2t+1}$. Thus the result will be proved by showing that $(4/d)^{2t}(d+1)^{2t+1}\leq d^{t}/et$.
Since ${\delta}\geq 27$ we have $d\geq 26$. Then it is easy to calculate that $f(d,t)=[4(1+1/d)]^{2t}(d+1)/etd^{t}<1$ when $d=26$ and $t=3$, and to observe that $f(d,t)$ decreases in $d$ and $t$. From this it follows that $(4/d)^{2t}(d+1)^{2t+1}\leq d^{t}/et$. ∎
At issue here is that it is not known if there exists a graph that satisfies the hypothesis of Theorem 3. Indeed, Biggs [3] defines a sequence of $k$-regular graphs $\{G_{i}\}$ with increasing $n(G_{i})$ to have large girth if ${\sf gir}(G_{i})\geq{\alpha}\log_{k-1}(n(G_{i}))$ for some constant ${\alpha}$. It is known that ${\alpha}\leq 2$, and the greatest known constant is a construction of [20] that yields ${\alpha}=4/3$. However, a graph satisfying the hypothesis of Theorem 3 necessarily has ${\alpha}=2$.
Fortunately, there are explicitly constructed girth 5 graphs that have $c(G)>{{\pi}^{*}}(G)$ (in fact, $c(G)\gg{{\pi}^{*}}(G)$), without referring to the conditions of Theorem 3. For examples of larger girth we still need Theorem 3.
@@
For all $d$ there exists a graph $G$ such that ${\sf gir}(G)=5$, ${{\pi}^{*}}(G)\leq 4$ and $c(G)\geq d$.
@@
Let $d$ be a large integer, and let $H$ be a $d$-regular graph with ${\sf gir}(H)\geq 5$. Create $G$ from $H$ by adding a leaf to each vertex of $H$, and then add an additional vertex $v$ adjacent to all of these leaves. Beginning with 4 pebbles at $v$, any vertex of $G$ can be reached in two pebbling steps, and so ${{\pi}^{*}}(G)\leq 4$. A robber, however, can choose to play only on the graph $H$. Because ${\sf gir}(G)=5$, it follows that if a cop occupies a vertex in $G$ different from the robber’s vertex, then this cop guards at most one of the robber’s neighbours. So if there are fewer than $d$ cops in $G$, the robber always has a safe neighbour to which to move and can evade capture indefinitely; i.e. $c(G)\geq d$. ∎
2.2 Upper Bounds
@@
Place two cops on each vertex of $S$ and one cop on each vertex of $S^{\prime}$. In order to not be immediately captured, the robber must start in $N[S]-S$, but then is captured in one step by some pair of cops from $S$. The second statement follows from choosing $S$ to be a minimum dominating set of $G$, since $S^{\prime}=\emptyset$. ∎
@@
For positive integers $m\geq 2k\geq 2$, let $Y=\{y_{1},\ldots,y_{m}\}$ and let $Q=\{Q_{1},\ldots,Q_{k}\}$ be a partition of $Y$ with each part size $|Q_{i}|\geq 2$. Define a bipartite graph $G$ with vertices $Y$, $Z=\{z_{1},\ldots,z_{k}\}$, and $x$ as follows. For each $1\leq j\leq k$ set $z_{j}\sim y_{i}$ if and only if $y_{i}\in Q_{j}$. Also set $x\sim y_{i}$ for every $1\leq i\leq m$. Then ${\gamma}(G)=k+1$. Indeed, since the neighborhoods of each $z_{j}$ are pairwise disjoint, at least $k$ vertices in $Y\cup Z$ are required to dominate $Z$, one from each $N[z_{j}]$. Suppose that $S$ is a dominating set of size $k$. By the above, $|S\cap N[z_{j}]|=1$ for all $j$. But to dominate $x$, some $y_{i}$ must be in $S$. Let $y_{i}\in N(z_{j})$; then $y_{i}$ does not dominate any other $y_{i^{\prime}}\in N(z_{j})$. Hence ${\gamma}(G)\geq k+1$. It is easy to see that $Z\cup\{x\}$ is a dominating set, so that ${\gamma}(G)=k+1$. By choosing $S=\{x\}$ we have $S^{\prime}=Z$, so that ${{\pi}^{\sf c}}(G)\leq k+2$, much better than $22{\gamma}(G)=2k+$ 2.
@@
Let $v$ be a vertex with ${\sf deg}(v)={\Delta}(G)$, set $S=\{v\}$, and apply Theorem 5 to obtain the general bound. Next, it is easy to see that ${{\pi}^{\sf c}}(K_{n})=n$ for $n\leq 2$. Then, a graph with at least three vertices has a vertex of degree at least two, so that $n-{\Delta}(G)+1\leq n-1$. Finally, if ${\Delta}(G)\geq 3$ then $n-{\Delta}(G)+1\leq n-2$, while if ${\Delta}(G)\leq 2$ then $G$ is a path or cycle, for which Theorem 18 yields ${{\pi}^{\sf c}}(G)=\lceil\frac{2n}{3}\rceil$, which is at most $n-2$ when $n\geq 6$. ∎
@@
The authors of [12] define a roman dominating set of $G$ to be a $\{0,1,2\}$-labeling of $V(G)$ so that every vertex labeled 00 is adjacent to some vertex labeled $2$. Note that the construction in Theorem 5 yields a roman dominating set by labeling each vertex by its number of cops. The roman domination number ${\gamma}_{R}(G)$ is defined to be the minimum sum of labels of a roman dominating set. Hence we obtain the following bound.
Theorem 7.
Every graph $G$ satisfies ${{\pi}^{\sf c}}(G)\leq{\gamma}_{R}(G)$.
@@
Suppose that ${{\pi}^{\sf c}}(H)\leq n(H)-s$. Then there is a configuration $C_{H}$ of $n(H)-s$ cops on $H$ that captures any robber on $H$. It remains to show that this number of cops can still win when the robber may move off $H$. Define the configuration $C_{G}$ of $n(G)-s$ cops on $G$ by placing one cop on each vertex of $G-H$ and $C_{H}(v)$ cops on each vertex $v\in H$. Then $C_{H}$ captures any robber on G. ∎
@@
One might be interested in measuring the gap between the size of a graph and its cop pebbling number. For this we define the cop deficiency of a graph $G$ to be Ï(${}^{\sf c}(G):=n(G)-{{\pi}^{\sf c}}(G)$. Then Theorem 8 and Corollary 9 can be restated as follows.
@@
For all $s\geq 2$ there is an $N=N(s)$ such that every graph $G$ with $n=n(G)\geq N$ has Ï(${}^{\sf c}(G)\geq s$.
@@
For example, let $T$ be a complete $k$-ary tree of depth $t$. Then ${\sf capt}(T)=t$ by Result 4, and so ${{\pi}^{\sf c}}(T)\leq 2^{t}$. Theorem 12 is tight for some graphs, as witnessed by any graph $G$ with a dominating vertex (see Corollary 16, below). It is also tight for any complete $k$-ary tree of depth two, when $k\geq 3$ (see Corollary 20, below).
@@
The result follows from Result 4 and Theorem 12. ∎
@@
Consider a maximum length path $P$ in $T$. Let $z$ be an endpoint of $P$ (necessarily a leaf), let $y$ be the neighbor of $z$, and let $x$ be the other neighbor of $y$ on $P$.
Base case: For $n=3$, the only tree is $P_{3}$. Place two cops on the central vertex, and the robber will be caught on the cops’ first move.
Inductive Step: Assume that for trees with $n<k$ vertices, ${{\pi}^{\sf c}}(T)\leq\lceil\frac{2n}{3}\rceil$. If $d(y)>2$, form a new tree $T^{\prime}=T-\{z\}-\{y\}-(\{N[y]-\{x\})\}$. By our inductive hypothesis, we can distribute the cops in such a way that the robber is caught if the robber starts on $T^{\prime}$. By placing two cops on $y$, we can also ensure that the robber is caught on the first move if the robber starts on $T-T^{\prime}$.
On the other hand, if $d(x)=d(y)=2$, form a new tree $T^{\prime}=T-\{x,y,z\}$. By our inductive hypothesis, we can distribute the cops in such a way that the robber is caught if the robber starts on $T^{\prime}$. By placing two cops on $y$, we can also ensure that the robber is caught if the robber starts on $T-T^{\prime}$.
If $d(y)=2$ and $d(x)>2$, and $x$ has a leaf neighbor $u$, form a new tree $T^{\prime}=T-\{u,y,z\}$. By our inductive hypothesis, we can distribute the cops in some distribution $D^{\prime}$ so that the robber is caught if the robber starts on $T^{\prime}$. By placing two cops on $y$, we can also ensure that the robber is caught if the robber starts on vertices $y$ or $z$. To capture a robber on $u$, one cop can reach $x$ from $D^{\prime}$, and another cop can reach $x$ from $y$. We then can reach $x$ from $u$.
Finally, suppose $d(y)=2$ and $d(x)>2$, and $x$ has no leaf neighbor $u$. Denote the neighborhood of $x$ which is not on $P$ as $N[x\cap P^{c}]=N_{2}[x]\cap T[V(T)\setminus V(P)]$, and let $u\in N[x]\cap T[V(T)\setminus V(P)]$. Since $P$ has maximum length, $N[u]-\{x\}$ consists only of leaves. Let $v\in N[u]-\{x\}$, and let $T^{\prime}=T-\{v,y,z\}$. By our inductive hypothesis, we can distribute the cops in some distribution $D^{\prime}$ that the robber is caught if the robber starts on $T^{\prime}$. If two cops can reach $x$ in $T^{\prime}$, we can add 2 more cops to $x$ to catch the robber on the vertices $\{v,y,z\}$. If two cops can reach $u$ in $T^{\prime}$, then $v$ and $x$ are reachable, so we can add 2 more cops to $y$ to catch the robber on the vertices $\{x,z\}$. Last, if two cops cannot reach $x$ or $u$, then no sequences of cop moves in $T^{\prime}$ will use the edge $uv$ (otherwise, we would be able to get two cops on at least one of the two vertices). Thus, we can simultaneously get one cop on $x$ and one cop on $u$. By adding two cops onto $y$, the cops can reach the vertices $\{v,y,z\}$. ∎
Note that, on a tree, cops move greedily toward the robber, so if a cop $p$ can reach a vertex $v$ then the robber cannot ever occupy $v$, as the robber has no access to $v$ except through $p$. Hence if $G$ is a tree then ${{\pi}^{\sf c}}(G)={{\pi}^{*}}(G)$. We note that Theorems 12 and 14 can each be stronger than each other, as the following two examples show. Define the spider $S(k,d)$ to be the tree having a unique vertex $x$ of degree greater than 2, all $k$ of whose leaves have distance $d$ from $x$.
@@
Let $J$ be such that $T\cap T_{j}\not=\emptyset$ if and only if $j\in J$, and set $Q_{j}=T\cap T_{j}$. Note that ${\sf gir}(G)\geq 4d-1$ implies that, for each $j\in J$, there is some $v\in T$ such that $Q_{j}\subseteq P(v)$. Moreover, by the definition of $S$, we have $\cup_{j\in J}Q_{j}=T$. In addition, ${\sf gir}(G)\geq 4d-1$ implies that, for each $j\in J$, the shortest $v_{i}v$-path $P^{*}_{i}$ is unique.
@@
For $1\leq i\leq 3$, define the tree $T_{i}$ to be the complete binary tree of depth $d-1$, rooted at vertex $v_{i}$, and define the tree $T$ to be the union of the three $T_{i}$ with an additional root vertex adjacent to each $v_{i}$. Then ${\gamma}_{d}(T)=d$, and $n=3(2^{d}-1)+1$, so that the bound from Theorem 15 is stronger than the bound from Theorem 14.

### v4

@@ 1 Introduction
We define these three paradigms more specifically in Subsection 1.1 below. The new graph invariant we define to study in this paper is the cop pebbling number of a graph, denoted ${{\pi}^{\sf c}}(G)$; roughly, this equals the minimum number of pebble-cops necessary to capture the robber in the Cops and Robbers Pebbling paradigm. In Subsections 1.21.4 we present known results about Cops and Robbers, dominating sets, and optimal pebbling, respectively, that will be used in the sequel. We record in Subsections 2.12.3 new theorems on lower bounds, upper bounds, and exact answers for ${{\pi}^{\sf c}}(G)$, respectively, for a range of graph families, including paths, cycles, trees, chordal graphs, high girth graphs, and cop-win graphs, as well as, in some cases, for all graphs. The reader already familiar with the prior subjects can skip ahead to Section 2 where we state and prove our results of the cop pebbling numbers of graphs. For example, Theorem 14 proves that the cop pebbling number of $n$-vertex trees is at most $\lceil 2n/3\rceil$, which is tight for paths (and cycles), and Theorem 15 provides an upper bound involving a domination parameter, as a function of girth. Section 3 contains theorems for Cartesian products of graphs. For example, Theorems 21 and 22 give upper and lower bounds on the cop pebbling numbers of grids and cubes, respectively, while Theorem 23 proves that ${{\pi}^{\sf c}}(G\Box K_{t})\leq t{{\pi}^{\sf c}}(G)$ for all $G$. Furthermore, Theorems 25 and 26 show that the analogous inequality from Graham’s Pebbling Conjecture fails for cop pebbling. We finish in Section 4 with some natural questions left open from this work, including a version of Meyniel’s Cops and Robbers Conjecture that may hold in the Cops and Robbers Pebbling world, namely that ${{\pi}^{\sf c}}(G)\leq 2n/3+o(n)$ for all $G$.
@@ 1.1 Definitions
In order to describe cop pebbling in Subsection 1.1.3, we need first to share the definitions and notations from each of the other two areas, as well as what we use from basic graph theory.
@@ 1.1.3 Cop Pebbling
The cop pebbling number ${{\pi}^{\sf c}}(G)$ is defined as the minimum number $m$ so that, from some configuration of $m$ cops, it is possible to capture any robber via pebbling steps. For example, $c(G)\leq n(G)$, since the placement of one cop on each vertex has already caught any robber. We call an instance of a graph $G$, configuration $C$, and robber vertex $v$ a game, and say that the cops win the game if they can capture the robber; else the robber wins. Note that, since we lose a cop with each pebbling step, the cops must catch the robber within at most |C|1|C|-1 turns to win the game.
@@ 1.2 Cop Results
Here we list the results on Cops and Robbers that will be used to prove our theorems on cop pebbling. A graph $G$ is cop-win if $c(G)=1$. A vertex $u$ in $G$ is called a corner if there is a vertex $v\not=u$ such that $N[u]\subseteq N[v]$. We say that $G$ is dismantlable if either $G$ is a single vertex or there is a corner $u$ such that GuG-u is dismantlable. Note that chordal graphs are dismantlable.
@@
Result 3 yields the following result as a corollary.
Result 4.
@@ 1.3 Dominating Set Results
Result 6 implies that ${\gamma}(Q^{d})\geq 2^{d}/(d+1)$. The following result shows that the actual value is not much greater, asymptotically.
Result 9.
@@ 1.4 Optimal Pebbling Results
The next two results involve Cartesian products.
Result 15.
@@
This bound is much smaller than the $\frac{2}{7}km+8\approx 0.2857km$ upper bound proved in [17], which is conjectured to be best possible.
Combining Results 11 and 12 yields the lower bound of the following result. The upper bound places $2^{k}$ pebbles on each vertex of a distance-$k$ dominating set, with $k$ roughly $d/3$.
Result 16.
@@ 2.1 Lower Bounds
The first two theorems give the cop number and optimal pebbling number as lower bounds to the cop pebbling number.
@@
If ${{\pi}^{\sf c}}(G)=c(G)$ then capturing the robber requires no steps. That is, if a pebbling capture requires a pebbling step, then the cop lost during the step is irrelevant in the Cops and Robbers capture, which contradicts the equality. That means that a successful pebbling configuration has no vertex without a pebble; i.e. ${{\pi}^{\sf c}}(G)=n(G)$. However, if $n(G)\geq 3$ then, by placing two cops at a vertex $v$ of degree at least two and one cop on each vertex not adjacent to $v$, then we can capture any robber in one step. (We especially highlight this statement, recorded as Corollary 7 in Subsection 2.2, below.) Thus ${{\pi}^{\sf c}}(G)<n$, a contradiction. If $n(G)=2$ then $G=K_{2}$ and ${{\pi}^{\sf c}}(K_{2})=2>1=c(K_{2})$, a contradiction. Hence $G=K_{1}$. ∎
@@
For any graph $G$, ${{\pi}^{\sf c}}(G)\geq{{\pi}^{*}}(G)$.
@@
Any configuration of $k$ cops, where $k<{{\pi}^{*}}(G)$, will contain a vertex $v$ which is unreachable. The robber can then choose to start and stay on $v$ and thus not be captured. ∎
@@ 2.2 Upper Bounds
The cycle $C_{n}$ satisfies ${{\pi}^{\sf c}}(C_{n})\leq\lceil\frac{2n}{3}\rceil$.
@@
For $C_{n}$, partition $C_{n}$ into $\lfloor\frac{n}{3}\rfloor$ copies of $P_{3}$ and, possibly, an extra $P_{1}$ or $P_{2}$. Place two cops on the center vertex of each $P_{3}$, and one cop on each vertex of the remaining one or two vertices. The robber can only choose to start on one of the copies of $P_{3}$, where he is next to a pair of cops, and so will be captured on the first move. Thus ${{\pi}^{\sf c}}(C_{n})\leq\lceil\frac{2n}{3}\rceil$. ∎
@@
If $G$ is a tree then ${{\pi}^{\sf c}}(G)\leq{{\pi}^{*}}(G)$.
@@
If $G$ is a tree then place ${{\pi}^{*}}(G)$ cops according to an optimal pebbling configuration $C$. The robber beginning at some vertex $v$ uniquely defines the subtree $T$ containing $v$ for which every cop in $T$ is on a leaf of $T$, and any leaf of $T$ with no cop is a leaf of $G$. Thus the robber can never escape $T$.
Now the cops follow the same greedy strategy used by optimal pebbling; i.e. at each state, they make all possible pebbling moves toward $v$. The new tree $T^{\prime}$ defined by the robber’s new vertex $v^{\prime}$ and resulting cop pebbles has strictly fewer vertices than $T$. Hence this process ends after finitely many steps.
Suppose that we allow, for the moment, that pebbles can be split into fractions, so that a fractional pebbling step moves half the amount of pebbles and removes the other half. Then the process above ends with the capture of the robber at some vertex $v^{*}$. Because the set of such pebbling steps is greedy with respect to each temporary robber position, they are greedy with respect to vertex $v^{*}$. That is, they are precisely the steps taken to solve $v^{*}$ from $C$ if the target $v^{*}$ was known in advance.
Therefore, because $C$ is an optimal pebbling configuration, each of the fractional steps are actually integral; i.e. $C$ captures the robber with pebbling steps. Hence ${{\pi}^{\sf c}}(G)\leq{{\pi}^{*}}(G)$, and the equality follows. ∎
@@
Place two cops on each vertex of $S$ and one cop on each vertex of $S^{\prime}$. In order to not be immediately captured, the robber must start in N[S]SN[S]-S, but then is captured in one step by some pair of cops from $S$. The second statement follows from choosing $S$ to be a minimum dominating set of $G$, since $S^{\prime}=\emptyset$. ∎
The authors of [12] define a Roman dominating set of $G$ to be a $\{0,1,2\}$-labeling of $V(G)$ so that every vertex labeled 0 is adjacent to some vertex labeled $2$. Note that the construction in Theorem 5 yields a Roman dominating set by labeling each vertex by its number of cops. The Roman domination number ${\gamma}_{R}(G)$ is defined to be the minimum sum of labels of a Roman dominating set. Hence we obtain the following bound.
Corollary 6.
Every graph $G$ satisfies ${{\pi}^{\sf c}}(G)\leq{\gamma}_{R}(G)$.
@@
For positive integers $m\geq 2k\geq 2$, let $Y=\{y_{1},\ldots,y_{m}\}$ and let $Q=\{Q_{1},\ldots,Q_{k}\}$ be a partition of $Y$ with each part size $|Q_{i}|\geq 2$. Define a bipartite graph $G$ with vertices $Y$, $Z=\{z_{1},\ldots,z_{k}\}$, and $x$ as follows. For each $1\leq j\leq k$ set $z_{j}\sim y_{i}$ if and only if $y_{i}\in Q_{j}$. Also set $x\sim y_{i}$ for every $1\leq i\leq m$. Then ${\gamma}(G)=k+1$. Indeed, since the neighborhoods of each $z_{j}$ are pairwise disjoint, at least $k$ vertices in $Y\cup Z$ are required to dominate $Z$, one from each $N[z_{j}]$. Suppose that $S$ is a dominating set of size $k$. By the above, $|S\cap N[z_{j}]|=1$ for all $j$. But to dominate $x$, some $y_{i}$ must be in $S$. Let $y_{i}\in N(z_{j})$; then $y_{i}$ does not dominate any other $y_{i^{\prime}}\in N(z_{j})$. Hence ${\gamma}(G)\geq k+1$. It is easy to see that $Z\cup\{x\}$ is a dominating set, so that ${\gamma}(G)=k+1$. By choosing $S=\{x\}$ we have $S^{\prime}=Z$, so that ${{\pi}^{\sf c}}(G)\leq k+2$, much better than $2{\gamma}(G)=2k+2$.
@@
Let $v$ be a vertex with ${\sf deg}(v)={\Delta}(G)$, set $S=\{v\}$, and apply Theorem 5 to obtain the general bound. Next, it is easy to see that ${{\pi}^{\sf c}}(K_{n})=n$ for $n\leq 2$. Then, a graph with at least three vertices has a vertex of degree at least two, so that $n-{\Delta}(G)+1\leq n-1$. Finally, if ${\Delta}(G)\geq 3$ then $n-{\Delta}(G)+1\leq n-2$, while if ${\Delta}(G)\leq 2$ then $G$ is a path or cycle, for which Theorem 18 yields ${{\pi}^{\sf c}}(G)=\lceil\frac{2n}{3}\rceil$, which is at most n2n-2 when $n\geq 6$. ∎
@@
Suppose that ${{\pi}^{\sf c}}(H)\leq n(H)-s$. Then there is a configuration $C_{H}$ of n(H)sn(H)-s cops on $H$ that captures any robber on $H$. It remains to show that this number of cops can still win when the robber may move off $H$. Define the configuration $C_{G}$ of n(G)sn(G)-s cops on $G$ by placing one cop on each vertex of GHG-H and $C_{H}(v)$ cops on each vertex $v\in H$. Then $C_{H}$ captures any robber on G. ∎
@@
One might be interested in measuring the gap between the size of a graph and its cop pebbling number. For this we define the cop deficiency of a graph $G$ to be $^{\sf c}(G):=n(G)-{{\pi}^{\sf c}}(G)$. Then Theorem 8 and Corollary 9 can be restated as follows.
@@
For all $s\geq 2$ there is an $N=N(s)$ such that every graph $G$ with $n=n(G)\geq N$ has $^{\sf c}(G)\geq s$.
@@
For example, let $T$ be a complete $k$-ary tree of depth $t$. Then ${\sf capt}(T)=t$ by Result 2, and so ${{\pi}^{\sf c}}(T)\leq 2^{t}$. Theorem 12 is tight for some graphs, as witnessed by any graph $G$ with a dominating vertex (see Corollary 16, below). It is also tight for any complete $k$-ary tree $T$ of depth two, when $k\geq 3$ (the lower bound follows from Theorem 18 since $T$ contains $P_{5}$).
@@
The result follows from Result 2 and Theorem 12. ∎
@@
This follows from Theorem 4 and Result 10. ∎
We note that Theorems 12 and 14 can each be stronger than each other, as the following two examples show. Define the spider $S(k,d)$ to be the tree having a unique vertex $x$ of degree greater than 2, all $k$ of whose leaves have distance $d$ from $x$.
@@
Let $J$ be such that $T\cap T_{j}\not$$\emptyset$ if and only if $j\in J$, and set $Q_{j}=T\cap T_{j}$. Note that ${\sf gir}(G)\geq 4d-1$ implies that, for each $j\in J$, there is some $v\in T$ such that $Q_{j}\subseteq P(v)$. Moreover, by the definition of $S$, we have $\cup_{j\in J}Q_{j}=T$. In addition, ${\sf gir}(G)\geq 4d-1$ implies that, for each $j\in J$, the shortest $v_{i}v$-path $P^{*}_{i}$ is unique.
@@
For $1\leq i\leq 3$, define the tree $T_{i}$ to be the complete binary tree of depth d1d-1, rooted at vertex $v_{i}$, and define the tree $T$ to be the union of the three $T_{i}$ with an additional root vertex adjacent to each $v_{i}$. Then ${\gamma}_{d}(T)=d$, and $n=3(2^{d}-1)+1$, so that the bound from Theorem 15 is stronger than the bound from Theorem 14.